OpenStudy (anonymous):

Find the derivative of f(x) = -3/x at x = -4.

OpenStudy (anonymous):

that's not an option, my options are 3/4 4/3 16/3 3/16

OpenStudy (anonymous):

oh i see, do you want to show how to get the answer (3/16)?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

Do you know the power rule?

OpenStudy (anonymous):

I know it but i'm confused how to use it

OpenStudy (anonymous):

and I thought with fractions, you couldnt use it

OpenStudy (anonymous):

first of we have -3/x we need to put the x in the numerator to use the power rule

OpenStudy (anonymous):

so multiply by -x on top and bottom?

OpenStudy (anonymous):

when we switch the x to the numerator, it will be x^-1 so it will be -3x^-1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

think of it as a reciprocal

OpenStudy (anonymous):

now we use the power rule, do you the formula?

OpenStudy (anonymous):

d/dx ^xn= nx^n-1

OpenStudy (anonymous):

yes! i learned better way if you use the f'(x)= cnx^n-1 where c= constant n= exponent

OpenStudy (anonymous):

its pretty much the same thing

OpenStudy (anonymous):

-3x^-1 use the power rule, we have -3(-1)x^-1-1

OpenStudy (anonymous):

then simplify that and sub in the -4 into x

OpenStudy (anonymous):

-4^-2 -4/x^2?

OpenStudy (anonymous):

-3(-1) multiply = 3x^-2

OpenStudy (anonymous):

oh, and then plug in -4? so: 3(-4)^2???

OpenStudy (anonymous):

yes, remember the the negative sign, change to the reciprocal it wil be 3(1/(-4^2))

OpenStudy (anonymous):

3(1/16)

OpenStudy (anonymous):

f'(x)= 3/16

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

glad i could help! :)